how many liters of 35% saline solution should be mixed with 65% saline solution to obtain 120 liters of 55% saline soluton?
Answer by stanbon(75887)

how many liters of 35% saline solution should be mixed with 65% saline solution to obtain 120 liters of 55% saline soluton?
-------
Equation:
salt + salt = salt
0.35x + 0.65(120-x) = 0.55*120
Multiply thru by 100 to get:
-----
35x + 65*120 - 65x = 55*120
-30x = -10*120
---
x = (1/3)(120)
x = 40 liters (amt. of 35% solution needed)
====================
Cheers,
Stan H.
================